# Simple and Provable Scaling Laws for the Test-Time Compute of Large Language Models

### v2

@@ Appendix B Examples for understanding and comparing the assumptions
$\mu_{A}$ $=p_{A}\times 0.5+p_{B}\times p_{\text{comp}}=0.5\times\alpha+p_{\text{comp}}\times(1-\alpha)=(0.5-p_{\text{comp}})\times\alpha+p_{\text{comp}},$
$\mu_{B}$ $=p_{B}\times 0.5+p_{A}\times(1-p_{\text{comp}})=0.5\times(1-\alpha)+(1-p_{\text{comp}})\times\alpha=(0.5-p_{\text{comp}})\times\alpha+0.5,$
which implies
@@
regardless of the value of $\alpha$.
@@
Appendix C Additional empirical results
C.1 Results for the knockout-style algorithm
@@
A comparison with majority voting.
Figure 12 shows that the knockout-style algorithm achieve higher accuracy than majority voting does in some cases, e.g., for the Qwen2.5 and Mixed options in GPQA, but not in all cases. Note that the knockout-style algorithm requires more test-time compute than majority voting does in this comparison; nonetheless, based on the trends shown in the figure, it is most likely that majority voting will not benefit much from further increasing its budget [3].
Figure 12: A comparison between the knockout-style algorithm (solid) and majority voting (dotted).
@@
Figure 13 includes empirical results of the knockout-style algorithm for each category of GPQA, while Figures 14 and 15 include those for MMLU-Pro-S.
@@
C.2 Results for the league-style algorithm
@@
Figure 16: The distribution of problems from the GPQA (top) or MMLU-Pro-S (bottom) benchmark, characterized by $\widehat{p}_{\text{comp}}$ from the knockout-style algorithm and $\widehat{\Delta}$ from the league-style algorithm (both with $N=16$) using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. The following abbreviations are used for the legend: K — knockout, L — league, R — right, W — wrong. For example, “KW-LR (# = 66)” means that there are 66 problems for which the knockout-style algorithm did wrong while the league-style algorithm did right.
@@
C.3 Prompt templates
Tables C.3 and C.3 include the prompt templates used in our experiments for generating a candidate solution and for comparing a pair of candidates, respectively. Some parts of our prompts, as well as code for parsing LLMs’ responses and extracting the answers for evaluation, are modified from those in the official GitHub repository of MMLU-Pro [^6] https://github.com/TIGER-AI-Lab/MMLU-Pro/tree/main.
@@ C.3 Prompt templates
% System prompt Please read the following multiple-choice questions and provide the most likely correct answer based on the options given. % User prompt # Question {question} # Output Format “‘ <reason>your step-by-step reasoning proecss</reason> <answer>"the answer is (X)" where X is the correct letter choice</answer> “‘
@@
% System prompt You are an impartial Judge. Given a question and two candidate solutions, your task is to choose which solution answer the question better. Your judgment should be unbiased, without favoring either Solution 1 or 2. % User prompt —- QUESTION —- {question} —- Solution 1 —- {candidate_a} —- Solution 2 —- {candidate_b} —- OUTPUT FORMAT —- “‘ <compare>compare both candidate solutions step-by-step thoroughly, and double check if there are mistakes in either solution</compare> <winner>Solution 1 or Solution 2 or Tie</winner> “‘

### v3

@@ Appendix B Examples for understanding and comparing the assumptions
$\mu_{A}$ $=p_{A}\times 0.5+p_{B}\times p_{\text{comp}}=0.5\times\alpha+p_{\text{comp}}\times(1-\alpha)$
$=(0.5-p_{\text{comp}})\times\alpha+p_{\text{comp}},$
$\mu_{B}$ $=p_{B}\times 0.5+p_{A}\times(1-p_{\text{comp}})=0.5\times(1-\alpha)+(1-p_{\text{comp}})\times\alpha$
$=(0.5-p_{\text{comp}})\times\alpha+0.5,$
which implies that
@@
is independent of the value of $\alpha$.
@@
Appendix C Supplementary materials for experiments
C.1 Additional implementation details
Throughout our experiments, the temperature for LLM decoding is set to 0.5 for the generation stage, and 0.1 for pairwise comparisons during the aggregation stage. Our early exploration suggests that these choices strike a good balance between diversity and preciseness in LLM decoding.
During the generation stage, we ask the LLM (via zero-shot CoT prompting [18]) to generate a reasoning process first and then its final answer. For each pairwise comparison during the aggregation stage, we also leverage zero-shot CoT prompting and ask the LLM to think step by step before deciding which solution in the pair is more plausible, unless specified otherwise. Tables C.3 and C.3 at the end of this section include the prompt templates used in our experiments for GPQA and MMLU-Pro-S, both of which are multiple-choice datasets. The prompt templates for MATH-500 are largely the same, only slightly adjusted to account for the desired output formats. Some parts of our prompts, as well as code for parsing LLMs’ responses and extracting the answers for evaluation, are modified from those in the official GitHub repository of MMLU-Pro [^4] https://github.com/TIGER-AI-Lab/MMLU-Pro/tree/main.
To account for the positional bias of LLMs [57, 41], we ensure that when a pair of candidates are compared for multiple times, they are placed in one order within the prompt for half of the comparisons, and in the opposite order for the other half.
C.2 Additional results for the knockout-style algorithm
Experiments with more datasets.
Figure 6 validates the efficacy of the knockout-style algorithm on MMLU-Pro-S and MATH-500.
Figure 6: Accuracy versus the number of initial candidates $N$ for the knockout-style algorithm on MMLU-Pro-S (left) and MATH-500 (right).
Distribution of problems.
Figure 7 illustrates the distribution of GPQA and MMLU-Pro-S problems, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ that are estimated with the empirical results for the knockout-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. Similarly, Figure 8 illustrates the results for QwQ-32B and GPT-4o.
Figure 7: The distribution of GPQA (top) and MMLU-Pro-S (bottom) problems, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ that are estimated with the empirical results for the knockout-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. Each plot is annotated with the number of problems satisfying the condition $\widehat{p}_{\text{comp}}>0.5$, $\widehat{p}_{\text{comp}}\leq 0.5$, $\widehat{p}_{\text{gen}}=0$ or $\widehat{p}_{\text{gen}}=1$. To the right of each plot is a histogram for $\widehat{p}_{\text{comp}}$. Each problem is represented by a circle if it is solved correctly by the knockout-style algorithm with $N=64$, and by a cross otherwise. We neglect problems with $\widehat{p}_{\text{gen}}=0$ or $1$, i.e., problems for which the initial candidate solutions are all incorrect or all correct, since there is no way of obtaining meaningful estimate of $\widehat{p}_{\text{comp}}$ for such problems.
Figure 8: The distribution of GPQA-diamond problems, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ that are estimated with the empirical results for the knockout-style algorithm using the QwQ-32B (left) or GPT-4o (right) option. Other settings are the same as those in Figure 7.
@@ C.2 Additional results for the knockout-style algorithm
Figure 11 includes empirical results of the knockout-style algorithm for each category of GPQA, while Figures 12 and 13 include those for MMLU-Pro-S.
@@
C.3 Additional results for the league-style algorithm
Results for MMLU-Pro-S.
Figure 14 includes empirical results for the league-style algorithm on MMLU-Pro-S.
Figure 14: Empirical results for the league-style algorithm on MMLU-Pro-S. Left: accuracy versus the number of initial candidates $N$ for the league-style (LG, solid lines) and knockout-style (KO, dotted lines) algorithms, given the same initial candidates. Right: accuracy versus $M$, the number of subsampled opponents for each candidate, for the league-style algorithm with $N=16$.
Distribution of problems.
Figure 15 illustrates the distribution of GPQA and MMLU-Pro-S problems, characterized by $\widehat{p}_{\text{cs}}$ and $\widehat{\Delta}$ that are estimated with the empirical results for the league-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option.
Figure 15: The distribution of GPQA (top) and MMLU-Pro-S (bottom) problems, characterized by $\widehat{p}_{\text{cs}}$ and $\widehat{\Delta}$ that are estimated with the empirical results for the league-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. Each plot is annotated with the number of problems satisfying the condition $\widehat{\Delta}>0$, $\widehat{\Delta}\leq 0$, $\widehat{p}_{\text{cs}}=0$ or $\widehat{p}_{\text{cs}}=1$. To the right of each plot is a histogram for $\widehat{\Delta}$. Each problem is represented by a circle if it is solved correctly by the league-style algorithm with $N=16$, and by a cross otherwise. We neglect problems with $\widehat{p}_{\text{cs}}=0$ or $1$, i.e., problems for which the initial candidate solutions are all incorrect or all correct, since there is no way of obtaining meaningful estimate of $\widehat{\Delta}$ for such problems.
@@ C.3 Additional results for the league-style algorithm
Figure 16: The distribution of GPQA (top) and MMLU-Pro-S (bottom) problems, characterized by $\widehat{p}_{\text{comp}}$ from the knockout-style algorithm and $\widehat{\Delta}$ from the league-style algorithm (both with $N=16$) using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. The following abbreviations are used for the legend: K — knockout, L — league, R — right, W — wrong. For example, “KW-LR (# = 66)” means that there are 66 problems for which the knockout-style algorithm did wrong while the league-style algorithm did right.
@@
Please read the following multiple-choice questions and provide the most likely correct answer based on the options given. # Question {question} # Output Format “‘ <reason>your step-by-step reasoning proecss</reason> <answer>"the answer is (X)" where X is the correct letter choice</answer> “‘
@@
You are an impartial Judge. Given a question and two candidate solutions, your task is to choose which solution answer the question better. Your judgment should be unbiased, without favoring either Solution 1 or 2. —- QUESTION —- {question} —- Solution 1 —- {candidate_a} —- Solution 2 —- {candidate_b} —- OUTPUT FORMAT —- “‘ <compare>compare both candidate solutions step-by-step thoroughly, and double check if there are mistakes in either solution</compare> <winner>Solution 1 or Solution 2 or Tie</winner> “‘